### v3

@@ Abstract
In this work, we have developed simple analytical tools of unprecedented tightness among EUR-based QKD security proofs [5]. Technically, the key enablers of these tools are a careful relaxation of the long-known additive Chernoff bound [20], and a classical link between random sampling with and without replacement [17]. Arguably, the tools might be of interest beyond the realm of QKD, in applications such as quality control tests, clinical trials or hypothesis testing in general. On top of it, we have put forward the possibility of replacing tail inequalities —the cornerstone of statistical analyses in QKD today— by exact computational methods in statistical inference. Remarkably, our results reveal the game-changing potential of such methods to achieve QKD with minimum block sizes, and reinforce the tightness of our simpler analytical tools when moderate-to-large blocks are considered.
@@
Acknowledgements. We acknowledge fruitful discussions with Jay Bartroff, Davide Rusca and Jean-Daniel Bancal. This work was supported by the Galician Regional Government (consolidation of Research Units: AtlantTIC), the Spanish Ministry of Economy and Competitiveness (MINECO), the Fondo Europeo de Desarrollo Regional (FEDER) through the grant No. PID2020-118178RB-C21, MICIN with funding from the European Union NextGenerationEU (PRTR-C17.I1) and the Galician Regional Government with own funding through the “Planes Complementarios de I+D+I con las Comunidades Autónomas” in Quantum Communication, the European Union’s Horizon Europe Framework Programme under the Marie Sklodowska-Curie Grant No. 101072637 (Project QSI) and the project QSNP (grant agreement No. 101114043). M.C. acknowledges support from a “Salvador de Madariaga” grant from the Spanish Ministry of Science, Innovation and Universities (grant No. PRX22/00192).
I Appendix I: ideal BBM92 protocol
I.1 Protocol description
@@ I.2 Failure probability
I.2.1 Relaxed additive Chernoff bound
@@
where $\Gamma^{+}_{n,\epsilon_{\rm PE}}(x)$ is given in Proposition 1 of the main text. This corresponds to the solid blue line of Fig. 2a there.
I.2.2 Serfling and Hush & Scovel inequalities combined
@@
As mentioned in the main text, this approach provides an optimized combination of Serfling inequality [14] with Hush and Scovel inequality [16], and it corresponds to the dashed orange line of Fig. 2a in the main text.
I.2.3 Serfling inequality
@@
This corresponds to the dotted green line of Fig. 2a in the main text.
I.2.4 Exact hypergeometric cumulative mass function
In the main text, a numerical confidence upper bound $\mathcal{CP}^{+}_{N,n,\epsilon}(\hat{p})$ on the population parameter is presented [27], based on the classical Clopper-Pearson method of pivoting the cumulative mass function [19]. Since one can replace $\Gamma^{+}_{n,\epsilon}(\hat{p})$ by $\mathcal{CP}^{+}_{N,n,\epsilon}(\hat{p})$ in Corollary 1 of the main text, $\mathcal{CP}^{+}_{N,n,\epsilon}(\hat{p})$ automatically fulfills Proposition 2 there as well. From this, one can easily show that it also satisfies Proposition 3, thereby providing another suitable threshold function:
@@ I.2.4 Exact hypergeometric cumulative mass function
This corresponds to the solid red line in of Fig. 2a in the main text.
II Appendix II: decoy-state BB84 protocol
@@ II.1 Protocol description
where $e_{1,\rm X}^{\rm th}=m_{1,\rm X}^{\rm th,U}/n_{1,\rm X}^{\rm th,L}$, =$\mathcal{B}=\bigl{\{}u\in\bigl{[}n_{1,\rm Z}^{\rm th,L}+n_{1,\rm X}^{\rm th,L},n_{1,\rm Z}^{\rm th,U}+n_{1,\rm X}^{\rm th,U}\bigr{]},v\in\bigl{[}n_{1,\rm X}^{\rm th,L},n_{1,\rm X}^{\rm th,U}\bigr{]},u>v\bigr{\}},$ and the threshold function under consideration is that of Eq. (13). Importantly, we adopt a finer-grained PE test for the random sampling tool of [15] in order to simplify the task of bounding the failure probability in Sec. II.3.2 below. Indeed, it is unclear whether a bound can be derived with this tool without modifying the PE test at all. After all, the tool in [15] is devised for the BBM92 protocol specifically, rather than for a decoy-state BB84 protocol, and in this work we provide a possible adaptation to the latter.
@@
    Parameter estimation (PE): For each $k\in\mathcal{K}$, they disclose the bit values in $\mathcal{X}_{k}^{\rm test}$ and compute the corresponding numbers of bit errors, $m_{{\rm X},k}$. With the available data, they perform the PE test that matches their preferred random sampling tool (see Sec. II.3 for the details). If the test fails, they abort the protocol.
@@
This approach yields provably $\left[2\left(\sqrt{\epsilon_{\rm PE}}+\delta\right)+\epsilon_{\rm PA}\right]$-secret [44, 43] and $\epsilon_{\rm cor}$-correct keys for the requested failure probability bound, $\epsilon_{\rm PE}$, carefully addressed in Sec. II.3.
@@
II.2 Decoy-state analysis
@@
where $\tau_{1}=\sum_{k\in\mathcal{K}}e^{-k}k\,p_{k}$ and $n_{\mathrm{Z},k}^{\pm}=B^{\pm}\left(\varepsilon,n_{{\rm Z},k},N_{\rm Z}\right)$ for the functions $B^{\pm}\left(\varepsilon,n_{{\rm Z},k},N_{\rm Z}\right)$ of Sec. II.2.1 to Sec. II.2.4, which depend on the tail inequality under consideration for Bernoulli sampling. Importantly as well, the superscript over the inequality symbol in Eq. (19) indicates that the corresponding bound holds except with probability $3\varepsilon$ at most. This follows from Boole’s inequality [47], noticing that all three statistical estimates have error probability $\varepsilon$ at most. We recall that this notation is extensively used below for convenience.
@@
$n_{1,\rm Z}\overset{2\varepsilon}{\leqslant}n^{\rm U}_{1,\rm Z}:=\tau_{1}\frac{n_{\rm Z,\nu}^{+}-n_{\rm Z,\omega}^{-}}{\nu-\omega}.$ (20)
@@
$m_{1,\rm X}\overset{2\varepsilon}{\leqslant}m^{\rm U}_{1,\rm X}:=\tau_{1}\frac{m_{\rm X,\nu}^{+}-m_{\rm X,\omega}^{-}}{\nu-\omega},$ (21)
@@
II.2.1 Relaxed additive Chernoff bound
First, we include the confidence intervals (CIs) that arise from our tight relaxation of the additive Chernoff bound [20]. Precisely, it follows from Proposition 1 in the main text that one can take $B^{\pm}(\varepsilon,n\hat{p},n):=n\Gamma^{\pm}_{n,\varepsilon}(\hat{p})$ for the statistics $\Gamma^{\pm}_{n,\varepsilon}(\hat{p})$ defined therein.
II.2.2 Exact multiplicative Chernoff bound
@@ II.2.3 Hoeffding inequality
II.2.4 Exact binomial cumulative mass function
Lastly, we include the optimal CIs presented in Appendix IV, due to [30]. These rely on the exact evaluation of the binomial cumulative mass function, which is determined by the regularized beta function. Precisely, one can take
@@ II.2.4 Exact binomial cumulative mass function
for any $\Delta>0$, where the statistics $\mathcal{F}^{\pm}_{n,\varepsilon}(\hat{p})$ are introduced in Theorem 1 of Appendix IV.
II.3 Failure probability
In this section, we bound the failure probability of the decoy-state BB84 protocol for the different random sampling tools under consideration.
II.3.1 Relaxed additive Chernoff bound
@@
where the threshold values $n_{1,\rm Z}^{\rm th}$ and $\phi_{1,\rm Z}^{\rm th}$ are the test-specific protocol inputs (see Sec. II.1), the random variable $n_{1,\rm Z}^{\rm L}$ is introduced in Sec. II.2, and the random variable $\phi_{1,\rm Z}^{\rm U}$ is defined as follows:
@@
where the variables $n_{1,\rm Z}^{\rm U}$, $n_{1,\rm X}^{\rm L(U)}$ and $m_{1,\rm X}^{\rm L(U)}$ are also presented in Sec. II.2. Besides, we recall that $\varepsilon>0$ is the common error probability settled for each individual concentration inequality in Sec. II.2.
@@
which follows because $A\implies{}B$ implies that $\Pr[A]\leq{}\Pr[B]$ for arbitrary events $A$ and $B$. Now, it is convenient to introduce the event where all the decoy-state bounds of Sec. II.2 hold. Namely,
@@
where we simply exploit the fact that $\Pr[A]\leq{}\Pr[\bar{B}]+\Pr[A,B]$ (for arbitrary events $A$ and $B$) and further invoke the trivial implication $\Lambda\implies n_{1,\rm Z}>n_{1,\rm Z}^{\rm L}$. Regarding the first term in the right-hand side of Eq. (33), carefully counting the errors in Sec. II.2 yields $\Pr\left[\bar{\Lambda}\right]\leq{}15\varepsilon$, and therefore we conclude
@@
follows. Note that this averaging technique is of the exact same nature as the one used to bound the failure probability of the BBM92 protocol (Proposition 3 in the main text), also deployed in [8, 15]. Lastly, combining this bound with Eqs. (36) and (34), we reach $\Pr\left[\Omega_{\rm TEST},\Omega_{\rm PE}\right]\leq{}16\varepsilon$. In conclusion, with the random sampling tool proposed in this work, $\epsilon_{\rm PE}=16\varepsilon$ provides the desired upper bound on the failure probability.
II.3.2 Serfling and Hush & Scovel inequalities combined
@@
where the threshold values $n_{1,\rm Z}^{\rm th,L(U)}$, $n_{1,\rm Z}^{\rm th,L(U)}$ and $m_{1,\rm X}^{\rm th,U}$ are the test-specific protocol inputs (see Sec. II.1). On the other hand, the event where the PE thresholds do not hold is again given by $\Omega_{\rm PE}=\{n_{1,\rm Z}\leq{}n_{1,\rm Z}^{\rm th}\}\cup\{\phi_{1,\rm Z}\geq{}\phi_{1,\rm Z}^{\rm th}\}$, recalling that, in this case, $n_{1,\rm Z}^{\rm th}$ and $\phi_{1,\rm Z}^{\rm th}$ are defined in Eq. (17) in terms of the thresholds above.
@@
which matches $\Lambda$ in Eq. (32) exactly except from the fact that it does not contemplate the condition $\{m_{1,\rm X}>m_{1,\rm X}^{\rm L}\}$. In particular, this implies that $\Pr[\bar{\Sigma}]=12\varepsilon$ (composing the relevant errors in Sec. II.3), and therefore $\Pr\left[\Omega_{\rm TEST},\Omega_{\rm PE}\right]\leq{}12\varepsilon+\Pr\left[\Omega_{\rm TEST},\Omega_{\rm PE},\Sigma\right]$. Crucially at this point, the ad hoc definitions of $\Omega_{\rm TEST}$ and $\Sigma$ assure the following trivial implications:
@@
where we have introduced the single-photon bit-error rate $e_{1,\rm X}=m_{1,\rm X}/n_{1,\rm X}$. Namely, the conjunction of $\Omega_{\rm TEST}$ and $\Sigma$ guarantees that the actual single-photon variables —and not only their decoy-state bounds— fulfil the thresholds of the PE test (first implication in Eq. (II.3.2)), which in turn implies that the random variable $e_{1,\rm X}=m_{1,\rm X}/n_{1,\rm X}$ is upper-bounded by the value $e_{1,\rm X}^{\rm th}=m^{\rm th,U}_{1,\rm X}/n^{\rm th,L}_{1,\rm X}$, and the value $\phi_{1,\rm Z}^{\rm th}=\max_{u,v,\in\mathcal{B}}q_{u,v,\epsilon_{\rm PE}}^{\rm th}(e_{1,\rm X}^{\rm th})$ is equal or larger than the function of random variables $q^{\rm th}_{\displaystyle{n_{1,\rm Z}+n_{1,\rm X},n_{1,\rm X},\varepsilon}}\left(e_{1,\rm X}^{\rm th}\right)$ (second implication in Eq. (II.3.2)). The latter holds simply because, if the actual single-photon RVs $n_{1,\rm Z}$ and $n_{1,\rm X}$ fulfil the thresholds of the PE test, the threshold function evaluated in these variables cannot be larger than the threshold function evaluated in the worst-case variables compatible with the thresholds (captivated by the maximization of some arbitrary inputs $u$ and $v$ over $\mathcal{B}$).
@@
where in the first inequality we invoke the first implication of Eq. (II.3.2) —and disregard the event $n_{1,\rm Z}\leq{}n_{1,\rm Z}^{\rm th}$ of $\Omega_{\rm PE}$ because it is incompatible with $n_{1,\rm Z}\in{}\bigl{(}n_{1,\rm Z}^{\rm th,L},n_{1,\rm Z}^{\rm th,U}\bigr{)}$—, in the second inequality we invoke the second implication of Eq. (II.3.2), and in the third inequality we make use of the fact that $\left\{\phi_{1,\rm Z}^{\rm th}\geq{}q^{\rm th}_{\displaystyle{n_{1,\rm Z}+n_{1,\rm X},n_{1,\rm X},\varepsilon}}\left(e_{1,\rm X}^{\rm th}\right),\ \phi_{1,\rm Z}\geq{}\ \phi_{1,\rm Z}^{\rm th}\right\}\implies{}\left\{\phi_{1,\rm Z}\geq{}\ q^{\rm th}_{\displaystyle{n_{1,\rm Z}+n_{1,\rm X},n_{1,\rm X},\varepsilon}}\left(e_{1,\rm X}^{\rm th}\right)\right\}$. To finish with, the desired bound follows from an averaging technique, as usual. Particularly, averaging over population sizes ($n_{1,\rm Z}+n_{1,\rm X}$) and test sample sizes ($n_{1,\rm X}$) in this case, it follows that
@@
II.3.3 Serfling inequality
With this tool [8, 14], we define the PE test of the protocol equally as in Sec. II.3.1, but taking
@@ II.3.3 Serfling inequality
Namely, we use the threshold function of Eq. (15) rather than that of Eq. (12). In these circumstances, one can readily prove the failure probability bound $\epsilon_{\rm PE}=13\varepsilon$ reproducing the steps of Sec. II.3.1, but with the ancillary event $\Sigma$ (Eq. (39)) rather than $\Lambda$ (Eq. (32)). This analogy with Sec. II.3.1 comes from the fact that the bound based on Serfling inequality can be trivially stated in the form of Proposition 2 in the main text, just like the bound proposed in this work.
II.3.4 Exact hypergeometric cumulative mass function
Similarly, the bound based on the Clopper-Pearson method [27, 19] can be stated in the form of Proposition 2 in the main text too. As a consequence, considering the simple PE test of Sec. II.3.1, one can establish that $\epsilon_{\rm PE}=16\epsilon$ following the exact same steps presented there —in particular, keeping the ancillary event $\Lambda$ of Eq. (32)—, but taking
@@ II.4 Channel model
Plugging the logarithmic inequality $\ln(1+z)\geq z(6+5z)/2(1+z)(3+z)$ [24] in each term of the divergence yields
@@
Let $\hat{p}$ be the average of $n$ independent Bernoulli variables, with expected value $\mathbb{E}\left(\hat{p}\right)=p$ such that $p\leq{}(1-2\kappa_{n,\epsilon})/(1+4\kappa_{n,\epsilon})$ for $\kappa_{n,\epsilon}=(2/9n)\ln(1/\epsilon)$. Then, $\Pr[\hat{p}\geq{}z_{n,\epsilon}(p)]\leq{\epsilon}$ with

### v4

@@ Abstract
In this work, we have developed statistical tools of unprecedented tightness among QKD security proofs [5]. Although we devise the tools for the BBM92 and the decoy-state BB84 protocols, they are of use in, e.g., measurement-device-independent QKD [40, 48] and twin-field QKD [49, 50] as well. Technically, the key enablers of our main analytical results are a classical link between random sampling with and without replacement [17], and a careful relaxation of the long-known additive Chernoff bound [25]. Arguably, the tools might be of interest beyond the realm of QKD, in applications such as quality control tests, clinical trials or hypothesis testing in general. To support this claim, in Appendices V and VI we thoroughly compare our relaxed Chernoff bound with existing hypergeometric bounds suited to the same purpose. What is more, for Hush and Scovel inequality and a tight Bernstein-type [51] inequality due to Greene and Wellner [52], we further reproduce our analytical method and obtain new analytical tools that outperform our Chernoff-based confidence bound if small enough block sizes are considered.
On top of it, we have put forward the possibility of replacing tail inequalities—the cornerstone of statistical analyses in QKD today—by exact computational methods in statistical inference. Remarkably, our results reveal the game-changing potential of such methods to achieve QKD with minimum block sizes, and reinforce the tightness of our simpler analytical tools when moderate-to-large blocks are considered.
@@
Acknowledgements—We acknowledge discussions with Jay Bartroff, Davide Rusca and Jean-Daniel Bancal, and useful feedback from Jon Wellner. This work was supported by the Galician Regional Government (consolidation of Research Units: AtlantTIC), the Spanish Ministry of Economy and Competitiveness (MINECO), the Fondo Europeo de Desarrollo Regional (FEDER) through the Grant No. PID2020-118178RB-C21, MICIN with funding from the European Union NextGenerationEU (PRTR-C17.I1) and the Galician Regional Government with own funding through the “Planes Complementarios de I+D+I con las Comunidades Autónomas” in Quantum Communication, the European Union’s Horizon Europe Framework Programme under the Marie Sklodowska-Curie Grant No. 101072637 (Project QSI) and the project Quantum Secure Networks Partnership (QSNP) (Grant No. 101114043). M.C. acknowledges support from a “Salvador de Madariaga” grant from the Spanish Ministry of Science, Innovation and Universities (Grant No. PRX22/00192).
Appendix I: ideal BBM92 protocol
.1 Protocol description
@@
.2.1 Relaxed additive Chernoff bound
@@
where $\Gamma^{+}_{n,\epsilon_{\rm PE}}(x)$ is given in Proposition 1 of the main text. This corresponds to the solid blue line of Fig. 2(a) there.
.2.2 Serfling and Hush & Scovel inequalities combined
@@
As mentioned in the main text, this approach provides a combination of Serfling inequality [14] with Hush & Scovel inequality [16] based on the union bound, and it corresponds to the dashed orange line of Fig. 2(a) in the main text.
.2.3 Serfling inequality
@@
This corresponds to the dotted green line of Fig. 2(a) in the main text.
.2.4 Exact hypergeometric cumulative mass function
In the main text, a numerical confidence upper bound $\mathcal{CP}^{+}_{N,n,\epsilon}(\hat{p})$ on the population parameter is presented [20, 21], based on the classical Clopper-Pearson method of pivoting the cumulative mass function [23]. Since one can replace $\Gamma^{+}_{n,\epsilon}(\hat{p})$ by $\mathcal{CP}^{+}_{N,n,\epsilon}(\hat{p})$ in Corollary 1 of the main text, $\mathcal{CP}^{+}_{N,n,\epsilon}(\hat{p})$ automatically fulfills Proposition 2 there as well. From this, one can easily show that it also satisfies Proposition 3, thereby providing another suitable threshold function:
@@
This corresponds to the solid red line in of Fig. 2(a) in the main text.
Appendix II: decoy-state BB84 protocol
@@
where $e_{1,\rm X}^{\rm th}=m_{1,\rm X}^{\rm th,U}/n_{1,\rm X}^{\rm th,L}$, =$\mathcal{B}=\bigl{\{}u\in\bigl{[}n_{1,\rm Z}^{\rm th,L}+n_{1,\rm X}^{\rm th,L},n_{1,\rm Z}^{\rm th,U}+n_{1,\rm X}^{\rm th,U}\bigr{]},v\in\bigl{[}n_{1,\rm X}^{\rm th,L},n_{1,\rm X}^{\rm th,U}\bigr{]},u>v\bigr{\}},$ and the threshold function under consideration is that of Eq. (13). Importantly, we adopt a finer-grained PE test for the random sampling tool of [15] in order to simplify the task of bounding the failure probability in Sec. .5.2 below. Indeed, it is unclear whether a bound can be derived with this tool without modifying the PE test at all. After all, the tool in [15] is devised for the BBM92 protocol specifically, rather than for a decoy-state BB84 protocol, and in this work we provide a possible adaptation to the latter.
@@
    Parameter estimation (PE): For each $k\in\mathcal{K}$, they disclose the bit values in $\mathcal{X}_{k}^{\rm test}$ and compute the corresponding numbers of bit errors, $m_{{\rm X},k}$. With the available data, they perform the PE test that matches their preferred random sampling tool (see Sec. .5 for the details). If the test fails, they abort the protocol.
@@
This approach yields provably $\left[2\left(\sqrt{\epsilon_{\rm PE}}+\delta\right)+\epsilon_{\rm PA}\right]$-secret [54, 53] and $\epsilon_{\rm cor}$-correct keys for the requested failure probability bound, $\epsilon_{\rm PE}$, carefully addressed in Sec. .5.
@@
.4 Decoy-state analysis
@@
where $\tau_{1}=\sum_{k\in\mathcal{K}}e^{-k}k\,p_{k}$ and $n_{\mathrm{Z},k}^{\pm}=B^{\pm}\left(\varepsilon,n_{{\rm Z},k},N_{\rm Z}\right)$ for the functions $B^{\pm}\left(\varepsilon,n_{{\rm Z},k},N_{\rm Z}\right)$ of Sec. .4.1 to Sec. .4.4, which depend on the tail inequality under consideration for Bernoulli sampling. Importantly as well, the superscript over the inequality symbol in Eq. (19) indicates that the corresponding bound holds except with probability $3\varepsilon$ at most. This follows from Boole’s inequality [57], noticing that all three statistical estimates have error probability $\varepsilon$ at most. We recall that this notation is extensively used below for convenience.
@@
$n_{1,\rm Z}\overset{2\varepsilon}{\leqslant}n^{\rm U}_{1,\rm Z}:=\frac{\tau_{1}}{{\nu-\omega}}\left(\frac{e^{\nu}}{p_{\nu}}n_{\rm Z,\nu}^{+}-\frac{e^{\omega}}{p_{\omega}}n_{\rm Z,\omega}^{-}\right).$ (20)
@@
$m_{1,\rm X}\overset{2\varepsilon}{\leqslant}m^{\rm U}_{1,\rm X}:=\frac{\tau_{1}}{{\nu-\omega}}\left(\frac{e^{\nu}}{p_{\nu}}m_{\rm X,\nu}^{+}-\frac{e^{\omega}}{p_{\omega}}m_{\rm X,\omega}^{-}\right),$ (21)
@@
.4.1 Relaxed additive Chernoff bound
First, we include the confidence bounds that arise from our tight relaxation of the additive Chernoff bound [25]. Precisely, it follows from Proposition 1 in the main text that one can take
$B^{\pm}(\varepsilon,n\hat{p},n):=n\Gamma^{\pm}_{n,\varepsilon}(\hat{p})$ (23)
for the statistics $\Gamma^{\pm}_{n,\varepsilon}(\hat{p})$ defined therein.
.4.2 Exact multiplicative Chernoff bound
@@
.4.4 Exact binomial cumulative mass function
Lastly, we include the optimal one-sided confidence intervals presented in Sec. Appendix IV: Optimal confidence bounds for the average parameter of a set of independent Bernoulli variables, due to [35]. These rely on the exact evaluation of the binomial cumulative mass function, which is determined by the regularized beta function. Precisely, one can take
@@
for any $\Delta>0$, where the statistics $\mathcal{F}^{\pm}_{n,\varepsilon}(\hat{p})$ are introduced in Theorem 1 of Sec. Appendix IV: Optimal confidence bounds for the average parameter of a set of independent Bernoulli variables.
.5 Failure probability
In this section, we provide the failure probability estimation of the decoy-state BB84 protocol for the different random sampling tools under consideration.
.5.1 Relaxed additive Chernoff bound
@@
where the threshold values $n_{1,\rm Z}^{\rm th}$ and $\phi_{1,\rm Z}^{\rm th}$ are the test-specific protocol inputs (see Sec. .3), the random variable $n_{1,\rm Z}^{\rm L}$ is introduced in Sec. .4, and the random variable $\phi_{1,\rm Z}^{\rm U}$ is defined as follows:
@@
where the variables $n_{1,\rm Z}^{\rm U}$, $n_{1,\rm X}^{\rm L(U)}$ and $m_{1,\rm X}^{\rm L(U)}$ are also presented in Sec. .4. Besides, we recall that $\varepsilon>0$ is the common error probability settled for each individual concentration inequality in Sec. .4.
@@
which follows because $A\implies{}B$ implies that $\Pr[A]\leq{}\Pr[B]$ for arbitrary events $A$ and $B$. Now, it is convenient to introduce the event where all the decoy-state bounds of Sec. .4 hold. Namely,
@@
where we simply exploit the fact that $\Pr[A]\leq{}\Pr[\bar{B}]+\Pr[A,B]$ (for arbitrary events $A$ and $B$) and further invoke the trivial implication $\Lambda\implies n_{1,\rm Z}>n_{1,\rm Z}^{\rm L}$. Regarding the first term in the right-hand side of Eq. (34), carefully counting the errors in Sec. .4 yields $\Pr\left[\bar{\Lambda}\right]\leq{}15\varepsilon$, and therefore we conclude
@@
follows. Note that this averaging technique is of the exact same nature as the one used to bound the failure probability of the BBM92 protocol (Proposition 3 in the main text), also deployed in [8, 15]. Lastly, combining this bound with Eq. (37) and Eq. (35), we reach $\Pr\left[\Omega_{\rm TEST},\Omega_{\rm PE}\right]\leq{}16\varepsilon$. In conclusion, with the random sampling tool proposed in this work, $\epsilon_{\rm PE}=16\varepsilon$ provides the desired upper bound on the failure probability.
.5.2 Serfling and Hush & Scovel inequalities combined
@@
where the threshold values $n_{1,\rm Z}^{\rm th,L(U)}$, $n_{1,\rm Z}^{\rm th,L(U)}$ and $m_{1,\rm X}^{\rm th,U}$ are the test-specific protocol inputs (see Sec. .3). On the other hand, the event where the PE thresholds do not hold is again given by $\Omega_{\rm PE}=\{n_{1,\rm Z}\leq{}n_{1,\rm Z}^{\rm th}\}\cup\{\phi_{1,\rm Z}\geq{}\phi_{1,\rm Z}^{\rm th}\}$, recalling that, in this case, $n_{1,\rm Z}^{\rm th}$ and $\phi_{1,\rm Z}^{\rm th}$ are defined in Eq. (17) in terms of the thresholds above.
@@
which matches $\Lambda$ in Eq. (33) exactly except from the fact that it does not contemplate the condition $\{m_{1,\rm X}>m_{1,\rm X}^{\rm L}\}$. In particular, this implies that $\Pr[\bar{\Sigma}]=12\varepsilon$ (composing the relevant errors in Sec. .5), and therefore $\Pr\left[\Omega_{\rm TEST},\Omega_{\rm PE}\right]\leq{}12\varepsilon+\Pr\left[\Omega_{\rm TEST},\Omega_{\rm PE},\Sigma\right]$. Crucially at this point, the ad hoc definitions of $\Omega_{\rm TEST}$ and $\Sigma$ assure the following trivial implications:
@@
where we have introduced the single-photon bit-error rate $e_{1,\rm X}=m_{1,\rm X}/n_{1,\rm X}$. Namely, the conjunction of $\Omega_{\rm TEST}$ and $\Sigma$ guarantees that the actual single-photon variables —and not only their decoy-state bounds— fulfill the thresholds of the PE test (first implication in Eq. (.5.2)), which in turn implies that the random variable $e_{1,\rm X}=m_{1,\rm X}/n_{1,\rm X}$ is upper-bounded by the value $e_{1,\rm X}^{\rm th}=m^{\rm th,U}_{1,\rm X}/n^{\rm th,L}_{1,\rm X}$, and the value $\phi_{1,\rm Z}^{\rm th}=\max_{u,v,\in\mathcal{B}}q_{u,v,\epsilon_{\rm PE}}^{\rm th}(e_{1,\rm X}^{\rm th})$ is equal or larger than the function of random variables $q^{\rm th}_{\displaystyle{n_{1,\rm Z}+n_{1,\rm X},n_{1,\rm X},\varepsilon}}\left(e_{1,\rm X}^{\rm th}\right)$ (second implication in Eq. (.5.2)). The latter holds simply because, if the actual single-photon random variables $n_{1,\rm Z}$ and $n_{1,\rm X}$ fulfill the thresholds of the PE test, the threshold function evaluated in these variables cannot be larger than the threshold function evaluated in the worst-case variables compatible with the thresholds (captivated by the maximization of some arbitrary inputs $u$ and $v$ over $\mathcal{B}$).
@@
where in the first inequality we invoke the first implication of Eq. (.5.2) —and disregard the event $n_{1,\rm Z}\leq{}n_{1,\rm Z}^{\rm th}$ of $\Omega_{\rm PE}$ because it is incompatible with $n_{1,\rm Z}\in{}\bigl{(}n_{1,\rm Z}^{\rm th,L},n_{1,\rm Z}^{\rm th,U}\bigr{)}$—, in the second inequality we invoke the second implication of Eq. (.5.2), and in the third inequality we make use of the fact that $\left\{\phi_{1,\rm Z}^{\rm th}\geq{}q^{\rm th}_{\displaystyle{n_{1,\rm Z}+n_{1,\rm X},n_{1,\rm X},\varepsilon}}\left(e_{1,\rm X}^{\rm th}\right),\ \phi_{1,\rm Z}\geq{}\ \phi_{1,\rm Z}^{\rm th}\right\}\implies{}\left\{\phi_{1,\rm Z}\geq{}\ q^{\rm th}_{\displaystyle{n_{1,\rm Z}+n_{1,\rm X},n_{1,\rm X},\varepsilon}}\left(e_{1,\rm X}^{\rm th}\right)\right\}$. To finish with, the desired bound follows from an averaging technique, as usual. Particularly, averaging over population sizes ($n_{1,\rm Z}+n_{1,\rm X}$) and test sample sizes ($n_{1,\rm X}$) in this case, it follows that
@@
.5.3 Serfling inequality
With this tool [8, 14], we define the PE test of the protocol equally as in Sec. .5.1, but taking
@@
Namely, we use the threshold function of Eq. (15) rather than that of Eq. (12). In these circumstances, one can readily prove the failure probability bound $\epsilon_{\rm PE}=13\varepsilon$ reproducing the steps of Sec. .5.1, but with the ancillary event $\Sigma$ (Eq. (40)) rather than $\Lambda$ (Eq. (33)). This analogy with Sec. .5.1 comes from the fact that the bound based on Serfling inequality can be trivially stated in the form of Proposition 2 in the main text, just like the bound proposed in this work.
.5.4 Exact hypergeometric cumulative mass function
Similarly, the bound based on the Clopper-Pearson method [20, 21] can be stated in the form of Proposition 2 in the main text too. As a consequence, considering the simple PE test of Sec. .5.1, one can establish that $\epsilon_{\rm PE}=16\epsilon$ is a valid choice following the exact same steps presented there —in particular, keeping the ancillary event $\Lambda$ of Eq. (33)—, but taking
@@
Therefore, it suffices to show that $D(z||p)\geq{}D(z,p)$ for all $z\in[p,1]$. For this purpose, we build on the logarithmic inequality [29]
$\ln(1+x)\geq{}\frac{x(6+5x)}{2(1+x)(3+x)}.$ (51)
Although this inequality is only established for $-1<x\leq{}0$ in [29], it actually holds for all $x>-1$, which is required in this proof. To see this, we note that the difference
$h(x)=\ln(1+x)-\frac{x(6+5x)}{2(1+x)(3+x)}$ (52)
is non-negative for all $x>-1$. This is shown by the fact that the derivative of $h(x)$ is given by
$h^{\prime}(x)=\frac{x^{3}}{(1+x)^{2}(3+x)^{2}},$ (53)
revealing that $h(x)$ reaches a global minimum at $x=0$, given by $h(0)=0$. Plugging the inequality in each term of the divergence yields
@@
Let $\hat{p}$ be the average of $n$ independent Bernoulli variables, with expected value $\mathbb{E}\left(\hat{p}\right)=p$, and let $\epsilon>0$. If $p\leq{}(1-2\kappa_{n,\epsilon})/(1+4\kappa_{n,\epsilon})$ for $\kappa_{n,\epsilon}=(2/9n)\ln(1/\epsilon)$, then $\Pr[\hat{p}\geq{}z_{n,\epsilon}(p)]\leq{\epsilon}$ with